Aceves, Pedro Instituto de Matemáticas Unidad Morelia, Morelia, México

Ahn, Eunkyung Kyungpook National University, Daegu, Korea

Akinola, Richard University of Bath, Bath, United Kingdom

AlZhour, Zeyad Zarqa Private University, Zarqa, Jordan

Aminu, Abdulhadi University of Birmingham, Birmingham, United Kingdom

Andjelic, Milica Center for Research on Optimization and Control, Aveiro, Portugal

Arav, Marina Georgia State University, Atlanta, GA

Arciniega-Nevarez, José Antonio Instituto de Matematicas, UNAM, Cuernavaca

Aricò, Antonio Dipartimento di Matematica - Universitá di Cagliari, Cagliari, Italy

Ariza, Ma. de Lourdes Instituto de Matematicas, UNAM, Cuernavaca

Arroyo, María José UAM-Iztapalapa, Mexico, City

Avalos Gaytan, Vanesa Universidad Autónoma de Nuevo León, Nuevo León, México

Avila-Romero, Julio César Instituto de Matematicas, UNAM, Cuernavaca

Bardsley, John University of Montana, Missoula, Montana

Barrett, Wayne Brigham Young University, Provo, Utah

Barría, José Santa Clara University, Santa Clara, CA

Baur, Ulrike Chemnitz University of Technology, Chemnitz, Germany

Beattie, Christopher Virginia Tech, Blacksburg, VA

Bella, Tom University of Connecticut, Storrs

Bengochea, Gabriel Universidad Autónoma de la Ciudad de México, México D.F.

Benner, Peter TU Chemnitz, Fakultät für Mathematik, Chemnitz, Germany

Berman, Abraham Technion, Haifa, Israel

Bernal-Vilchis, Fernando Instituto de Matematicas, UNAM, Morelia

Bianchini, Barbara Pontificia Universidade Catolica de Sao Paulo, Sao Paulo

Bini, Dario Dipartimento di Matematica, Universitá di Pisa, Pisa, Italy

Boettcher, Albrecht TU Chemnitz, Chemnitz, Germany

Boimond, Jean-Louis LISA - University of Angers, Angers, France

Bourgeois, Gerald Faculté de Luminy, Marseille, France

Brualdi, Richard A. University of Wisconsin - Madison, Madison

Bru, Rafael Univ. Politécnica, Valencia, Spain

Bueno Cachadina, María Isabel The University of California at Santa Barbara, Santa Barbara

Bunch, James UCSD, La Jolla, CA

Butkovic, Peter University of Birmingham, Birmingham, United Kingdom

Canogar, Roberto Universidad Nacional de Educación a Distancia, Madrid, Spain

Carreón-Otañez, Cesar Fac. Ciencias, UNAM, Mexico city

Carriegos, Miguel Universidad de León, León, Spain

Castro-González, Nieves Universidad Politécnica de Madrid, Madrid, Spain

Catral, Minerva University of Victoria, Victoria, Canada

Cheng, Wei National University of Defense Technology, Changsha, P.R. China

Coba-Magaña, Alejandro Waldemar Instituto de Matematicas, UNAM, Morelia

Coronado-Juárez, Erick Instituto de Matemáticas, UNAM, Mexico city

Corral, Cristina Universidad Politécnica de Valencia, Valencia, Spain

Corriveau, Claudia Université du Québec à Montréal (UQAM), Montréal, Canada

Cortes-Aguirre, Javier de Jesus Fac. Ciencias, UNAM, Mexico city

Cortés, Vanesa Universidad de Zaragoza, Zaragoza, Spain

Costa, Liliana Centre for Research on Optimization and Control, Aveiro, Portugal

Cox, Steven Rice University, Houston, TX

Cravo, Glória University of Madeira and CELC, Funchal, Portugal

Cvetkovic, Ljiljana Faculty of Science, Novi Sad, Serbia

da Cruz, Henrique F. U.B.I, Covilhã, Portugal

Damm, Tobias TU Kaiserslautern, Kaiserslautern, Germany

Davila, Adriana Instituto de Matemáticas, UNAM, Mexico city

Day, Jane San Jose State University, San Jose, CA

Díaz-Gutiérrez, Alejandro Instituto de Matematicas, UNAM, Cuernavaca

Díaz-Gutiérrez, Ricardo Instituto de Matematicas, UNAM, Cuernavaca

Deaett, Louis University of Wisconsin-Madison, Madison, WI

DeAlba, Luz Drake University, Des Moines, IA

Deb, Biswajit Sikkim Manipal Institute of Technology, Majitar, Sikkim, India

Dhillon, Inderjit University of Texas, Austin

Djordjevic, Dragan Faculty of Sciences and Mathematics, University of Nis, Nis , Serbia

Dobovisek, Mirko University of Ljubljana, Faculty of Mathematics & Physics, Ljubljana, Slovenia

Dodig, Marija CELC, Universidade de Lisboa, Lisbon, Portugal

Dogan-Dunlap, Hamide UTEP, El Paso, TX

Dolinar, Gregor Faculty of Electrical Engineering, Ljubljana, Slovenia

Domínguez, María Elena Universidad Politécnica de Madrid, Madrid, Spain

Dopazo, Esther Facultad de Informática. Universidad Politécnica de Madrid, Boadilla del Monte, Madrid, Spain

Engel, Gernot Transversal Networks Inc., Madison Wisconsin

Vallejo, Ernesto Instituto de Matemáticas, UNAM, Morelia

Esen, Özlem Anadolu University, Eskisehir, Turkey

Estatico, Claudio Universitá di Cagliari, Cagliari, Italy

Eubanks, Sherod Washington State University, Pullman

Fassbender, Heike TU Braunschweig, Braunschweig, Germany

Feng, Lihong Faculty of Mathematics, TU Chemnitz, Chemnitz, Germany

Fernandes, Rosário Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal

Ferrer, Josep Universitat Politecnica de Catalunya, Barcelona, Spain

Fonseca, Carlos Department of Mathematics, University of Coimbra, Coimbra, Portugal

Fošner, Ajda Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

Frank, Martin University of Kaiserslautern, Kaiserslautern, Germany

Freund, Roland University of California, Davis, Davis, CA

Furtado, Susana Centro de Estruturas Lineares e Combinatórias da U. L, Porto, Portugal

Furuichi, Shigeru Nihon University, Tokyo, Japan

Gassó, Maria Teresa Inst. Mat. Mult. Universidad Politécnica Valencia, Valencia, Spain

Gaubert, Stephane INRIA and CMAP, Ecole Polytechnique, Palaiseau, France

Gavalec, Martin University of Hradec Králové, Hradec Králové, Czech Republic

Gemignani, Luca Department of Mathematics University of Pisa, Pisa, Italy

Glebsky, Lev Universidad Autónoma de San Luis Potosi, San Luis Potosi, México

Goldberger, Assaf Tel Aviv University, Tel Aviv, Israel

Gouveia, María Department of Mathematics, FCTUC, Coimbra, Portugal

Grout, Jason Iowa State University, Ames, IA

Grudsky, Sergey CINVESTAV del I.P.N., México, México

Gugercin, Serkan Virginia Tech, Blacksburg, VA

Guo, Chun-Hua University of Regina, Regina, Canada

Guzmán, José Ramón Instituto de Investigaciones Económicas. UNAM., México, México

Guzmán Ovando, María Esperanza Benemerita Universidad Autonoma de Puebla, Puebla

Hardouin, Laurent LISA - University of Angers, Angers, France

Harel, Guershon Dept of Math, University of California, San Diego, San Diego

Hernandez Escamilla, Esteban Librado UNAM, Mexico, D.F.

Hernández-Cedillo, Magdalena Instituto de Matematicas, UNAM, Mexico city

Herrera, Crispin ESFM-IPN, Mexico city

Hershkowitz, Daniel Technion, Haifa, Israel

Hnětynková, Iveta, Dep. of Mathematics, Arizona State University, Tempe Hogben, Leslie

Iowa State University, Ames, Iowa

Horn, Roger University of Utah, Salt Lake City

Iannazzo, Bruno Dipto. di Fisica e Mat., Università dell’Insubria, Como, Italy

Im, Bokhee Chonnam National University, Gwangju, Korea

Jiang, Er-Xiong Shanghai University, Shanghai, China

Jiménez-Munguía, Ronald Richard Instituto de Matematicas, UNAM, Cuernavaca

Jorge, Otero BUAP, Puebla

Kantun Montiel, Gabriel Benemérita Universidad Autónoma de Puebla, Puebla, México

Karow, Michael TU Berlin, Berlin, Germany

Kú, Darly CINVESTAV, Distrito Federal

Kirkland, Steve University of Regina, Regina, Canada

Klasa-Bompoint, Jacqueline Dawson College, Montreal, Canada

Klein, Andre University of Amsterdam, Amsterdam, The Netherlands

Kopparty, Bhaskara Rao Indiana State University, Terre Haute

Körtesi, Peter University of Miskolc, Miskolc, Hungary

Kostici, Vladimir Faculty of Science, Novi Sad, Serbia

Kressner, Daniel ETH Zurich, Zurich, Switzerland

Kum, Sangho Chungbuk National University, Cheongju, S. Korea

Laffey, Thomas University College Dublin, Dublin, Ireland

Lancaster, Peter University of Calgary, Calgary, Canada

Larson, Christine San Diego State University & Indiana University, San Diego

Lay, David University of Maryland, College Park, Maryland

Lee, Gue Myung Pukyong National University, Busan, Korea

Lee, Hosoo Kyungpook National University, Daegu, Korea

Leon, Steven University of Massachusetts Dartmouth, Dartmouth, MA

Li, Chi-Kwong College of William and Mary, Williamsburg, VA

Lim, Yongdo Kyungpook National University, Daegu, Korea

Loewy, Raphael Technion-Israel Institute of Technology, Haifa, Israel

Loiseau, Jean Jacques IRCCyN-CNRS, Nantes, France

López-Chávez, Gloria Instituto de Matematicas, UNAM, Mexico city

López-Trujillo, Alma Delia Instituto de Matematicas, UNAM, Cuernavaca

Luviano Flores, Johana Instituto de Matemáticas, UNAM, Mexico city

M. Dopico, Froilán Universidad Carlos III de Madrid, Leganés, Spain

Machado, Silvia Pontificia Universidade Catolica de São Paulo, São Paulo, Brasil

Maracci, Mirko Dept of Mathematics and CSCI, Siena University, Siena, Italy

Maranhao, Maria Cristina Pontificia Universidade Católica de São Paulo, São Paulo

Marovt, Janko Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

Marques, Maria-da-Graça University of Algarve and CELC, Faro, Portugal

Martinez-Avendano, Ruben Universidad Autonoma del Estado de Hidalgo, Pachuca

Martínez, José-Javier Universidad de Alcalá, Alcalá de Henares, Spain

Martin, William North Dakota State University, Fargo, USA

Martins, Enide Centre for Research on Optimization and Control (CEOC), Aveiro, Portugal

Martínez, Gerardo Escuela de Ciencias UABJO, Oaxaca

Mathewkutty, Habel Habel Math, Humble, TX

McDonald, Judith Washington State University, Pullman, WA

McEneaney, William University of California, San Diego, La Jolla

Mead, Jodi Boise State University, Boise

Meerbergen, Karl K.U.Leuven, Heverlee, Belgium

Meini, Beatrice Dipartimento di Matematica, Universitá di Pisa, Pisa, Italy

Mena, Hermann Escuela Politécnica Nacional, Quito, Ecuador

Merlet, Glenn CNRS/LIAFA, Paris, France

Merris, Russ None–Retired, Hayward, CA

Miguel-Ortiz, Daniel ESFM-IPN, Mexico city

Mikkelson, Rana Iowa State University, Ames, IA

Milligan, Thomas University of Central Oklahoma, Edmond, OK

Mishra, Pramod Kumar BHU Varanasi ( India), Varanasi

Mitchell, Lon Virginia Commonwealth University, Richmond

Moosaei, Hossein Guilan, Rasht, Iran

Moro, Julio Universidad Carlos III de Madrid, Leganés, Spain

Morris, DeAnne Washington State University, Pullman

Moura, Ana Instituto Superior Técnico, UTL, Lisbon, Portugal

Nagy, James Emory University, Atlanta, GA

Narayan, Sivaram Central Michigan University, Mount Pleasant, Michigan

Nava-Lara, Joel ESFM-IPN, Mexico city

Neumann-Coto, Max Instituto de Matemáticas, UNAM, Mexico city

Neumann, Michael University of Connecticut, Storrs

Olesky, Dale University of Victoria, Victoria, Canada

Olshevsky, Vadim University of Connecticut, Storrs

Palma, Alejandro Instituto de Física (BUAP), Puebla, México

Parraguez, Marcela Pontificia Universidad Católica de Valparaíso, Valparaíso, Chile

Patricio, Pedro Departamento de Matemática, Universidade do Minho, Braga, Portugal

Peña, Juan Manuel University of Zaragoza, Zaragoza, Spain

Peiña, Marta Universitat Politecnica de Catalunya, Barcelona, Spain

Perdigão, Cecília Faculdade de Ciências e Tecnologia-UNL, Lisboa, Portugal

Perez, Diego UNAM, Mexico

Plavka, Ján Technical University, Kocice, Slovakia

Plestenjak, Bor University of Ljubljana, Ljubljana, Slovenia

Ponce, Daniela University of Hradec Králové, Hradec Králové, Czech Republic

Poole, George East Tennessee State University, Johnson City

Possani, Edgar ITAM- Instituto Tecnológico Autónomo de México, México, México

Pérez-Tiscareño, Reyna María Instituto de Matematicas, UNAM, Mexico city

Prieto-Moreno, Kernel Enrique IIMAS UNAM, Mexico city

Prokip, Volodymyr Institute of Appl. Problem for Mech. and Math. NAS of Ukr, L’viv, Ukraine

Protasov, Vladimir Moscow State University, Moscow, Russia

Pruneda, Rosa E. University of Castilla-La Mancha, Ciudad Real, Spain

Pryporova, Olga Iowa State University, Ames, IA

Renaut, Rosemary Department of Mathematics and Statistics, Tempe

Rivera Martínez, Luis Manuel IICO, Universidad Autónoma de San Luis Potosí, San Luis Potosí

Roa Fuentes, Solange CINVESTAV, México Distrito Federal

Roca, Alicia Dpto. de Matemática Aplicada, Universidad Politécnica, Valencia, Spain

Rodríguez, Juan Alberto Universitat Rovira i Virgili, Tarragona, España

Rosenthal, Peter University of Toronto, Canada

Ross, Robert Elsevier, Cambridge, MA

Rump, Siegfried M. Hamburg University of Technology, Hamburg, Germany

Russo, Maria Rosaria Department of Mathematics - University of Padua, Padova, Italy

Rust, Bert National Institute of Standards and Technology, Gaithersburg, MD

Salam, Ahmed Université du Littoral-Côte d’Opale, Calais, France

Sanchez, Antonio Jesus ESFM-IPN, Mexico city

Sánchez Perales, Salvador Benemérita Universidad Autónoma de Puebla, Puebla, México

Santos, Pedro A. Instituto Superior Técnico, UTL, Lisbon, Instituto Superior Técnico, UTL, Lisbon, Portugal

Sarmiento-Rosales, Eliseo ESFM-IPN, Mexico city

Satô, Kenzi Tamagawa University, Tokyo, Japan

Schaeffer, Elisa Universidad Autónoma de Nuevo León, San Nicolás de los Garza, México

Schaffrin, Burkhard Ohio State University, Columbus, Ohio

Schneider, Hans University of Wisconsin, Madison, Madison, WI

Sebeldin, Anatoly University UGANC, Guinea, Conakry, Guinea

Seddighin, Morteza Indiana University East, Richmod

Semrl, Peter University of Ljubljana, Ljubljana, Slovenia

Sendov, Hristo The University of Western Ontario, London, Canada

Sergeev, Sergey University of Birmingham, Birmingham, United Kingdom

Shader, Bryan University of Wyoming, Laramie

Shahryari, Mohammad Tabriz University, Tabriz, Tabriz, Iran

Shaked-Monderer, Naomi Emek Yezreel College, Emek Yezreel, Israel

Shivakumar, Pappur University of Manitoba, Winnipeg, Canada

Singer, Ivan Romanian Academy of Sciences, Bucharest, Romania

Sinkovic, John Technische Universiteit Eindhoven, Eindhoven, Netherlands

Sivic, Klemen Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

Šmigoc, Helena University College Dublin, Dublin, Ireland

Soares, Graça, University of Tras-os-Montes and Alto Douro, Vila Real, Portugal Spitkovsky, Ilya

College of William and Mary, Blacksburg, VA, USA

Stefan, Wolfgang Arizona State University, Tempe

Stosic, Marko Instituto de Sistemas e Robotica, IST, Lisbon, Portugal

Strong, David Pepperdine University, Malibu, California

Stuart, Jeff Pacific Lutheran University, Tacoma, WA, USA

Szyld, Daniel Temple University, Philadelphia

Takane-Imay, Martha Instituto de Matemáticas UNAM (Cuernavaca), Cuernavaca

Tam, Bit-Shun Tamkang University, Tamsui, Taiwan

Tam, Kin Po University of Birmingham, Birmingham, United Kingdom

Tanguay, Denis Université du Québec à Montréal (UQAM), Montréal, Canada

Teixeira Matos, Isabel Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal

Treviño, Francisco universidad autónoma metropolitana Azcapotzalco, México

Trigueros, María Depto de Matemáticas ITAM, México DF

Uchiyama, Mitsuru Shimane University, Matsue, Shimane, Japan

Uhlig, Frank Mathematics, Auburn University, Auburn, AL

Vallejo, Ernesto Instituto de Matemáticas, Morelia, México

van den Driessche, Pauline University of Victoria, Victoria, Canada

van der Holst, Hein Eindhoven University of Technology, Eindhoven, The Netherlands

Vander Meulen, Kevin Redeemer University College, Ancaster, Ontario, Canada

Van Dooren, Paul Université catholique de Louvain, Louvain-la-Neuve, Belgium

Vargas, Xaab Nop ICYTDF, México, México

Velazquez-Guerrero, Luis Carlos Facultad de Ciencias, UNAM, Mexico city

Verde-Star, Luis UAM Iztapalapa, Mexico DF

Vieira, Luis Feup, Porto, Portugal

Villalobos-Sánchez, Armando Esteban Facultad de Ciencias, UNAM, Mexico city

Weaver, James University of West Florida, Pensacola, FL

Wojciechowski, Piotr University of Texas at El Paso, El Paso

Wrobel, Iwona Warsaw Univ. of Technology and Polish Academy of Sciences, Warsaw, Poland

Wu, Pei Yuan National Chiao Tung University, Hsinchu, Taiwan

Zandieh, Michelle Arizona State University, Tempe, AZ

Zimmermann, Karel Faculty of Mathematics and Physics, Charles University, Prague, Czech Republic

Zúñiga, Juan Carlos University of Calgary, Calgary, Canada